A warehouse manager at Mary Beth Marrs Corp. needs to simulate the demand placed on a product that does not fit standard models. The concept being measured is​ "demand during lead​ time," where both lead time and daily demand are variable. The historical record for this​ product, along with the cumulative​distribution, appear in the table.

                                                                                              

Demand During Lead Time

Probability

Cumulaive Probability

100

0.01

0.01

120

0.15

0.16

140

0.25

0.41

160

0.20

0.61

180

0.04

0.65

200

0.08

0.73

220

0.27

1.00

The following random numbers have been​ generated:

64​,

83​,

79​,

1​,

and

56.

​(Note: Assume the random number interval begins at 01 and ends at

00.​)

Based on the given probabilty​ distribution, for the given random number the demand during the lead time​ is:

Random Number

Demand

64

nothing

83

nothing

79

nothing

1

nothing

56

nothing

The average demand during the lead time is

nothing

​(enter your response as an​ integer).The total demand during the lead time based on the five simulations is

nothing

​(enter your response as an​ integer).

Enter your answer in each of the answer boxes.
Answer #1

Simulation of demand during lead time is following:

EXCEL FORMULAS:

Demand During Lead Time Probability Random number interval
100 0.01 0 =B2*100+C2
120 0.15 =D2 =B3*100+C3
140 0.25 =D3 =B4*100+C4
160 0.2 =D4 =B5*100+C5
180 0.04 =D5 =B6*100+C6
200 0.08 =D6 =B7*100+C7
220 0.27 =D7 00
Random Number Demand
64 =LOOKUP(A12,$C$2:$C$8,$A$2:$A$8)
83 =LOOKUP(A13,$C$2:$C$8,$A$2:$A$8)
79 =LOOKUP(A14,$C$2:$C$8,$A$2:$A$8)
1 =LOOKUP(A15,$C$2:$C$8,$A$2:$A$8)
56 =LOOKUP(A16,$C$2:$C$8,$A$2:$A$8)
The average demand during lead time is = =AVERAGE(B12:B16)